TREX: Logbook Entries

TREX: Site south Messages, 25 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
267 Mon 01-May-2006South data copiedsouthpoulos
259 Sat 29-Apr-2006Guy Wire Tensionssouthgolubieski
247 Wed 26-Apr-2006Trime Measurements, Gravimetricsouthagee
227 Sat 22-Apr-2006Guy wire tensionssouthpoulos
226 Sat 22-Apr-20065m, 30m Kryptons cleanedsouthmilitzer
217 Thu 20-Apr-2006Trime Measurement, Gravimetricsouthmilitzer
216 Thu 20-Apr-2006Rad.cleanedsouthmilitzer
198 Sat 15-Apr-2006Boom anglessouthpoulos
184 Thu 13-Apr-2006Updates on Echo/Qsoil negative at southsouthhorst
164 Sun 09-Apr-2006Radiometers 'Cleaned'southmilitzer
163 Sun 09-Apr-2006gravimetric soil moisturesouthmilitzer
146 Tue 04-Apr-2006loggers at West and Southsouthsemmer
144 Mon 03-Apr-2006Intermittent logger datasouthhorst
143 Mon 03-Apr-2006replaced MaxStream radio in logger at southsouthhorst
126 Thu 30-Mar-2006Qsoil negative at southsouthhorst
96 Thu 23-Mar-2006Maintenance at south yesterdaysouthhorst
63 Thu 16-Mar-2006South visitsouthoncley
55 Wed 15-Mar-2006Service visit to South towersouthhorst
52 Tue 14-Mar-200615 m TRH failed at south sitesouthhorst
49 Tue 14-Mar-2006South service visitsouthhorst
44 Sun 12-Mar-2006south visitsouthagee
27 Sun 05-Mar-2006south 30m now serializedsouthoncley
25 Sun 05-Mar-2006Serializer WARNING!southoncley
15 Tue 28-Feb-2006usb drive installsouthoncley
7 Mon 27-Feb-2006South network tasksouthoncley

267: data-system, Site south, Mon 01-May-2006 12:42:50 PDT, South data copied
The south pocketec data was copied to /media/exthd7 in the 'south' subdirectory. This is the final download of south data.
259: daily status, Site south, Sat 29-Apr-2006 17:04:21 PDT, Guy Wire Tensions
South 30M tower guy wire tensions.

Inner to outer
N:  325,275,250,280
SE: 360,310,350,300
SW: 375,275,340,290

247: soil, Site south, Wed 26-Apr-2006 14:59:36 PDT, Trime Measurements, Gravimetric
25 April, 2006
South Site

Sample taken at 1220PST
Samples placed in oven at 1300PST

0-3 cm  154.8-48.7 = 106.1g net

		Wet	Dry

	D1	30.5	30.2
	D2	33.7	33.5
	D3	41.9	41.5

3-6 cm 143.0-29.6 = 113.4 net

		Wet	Dry

	E1	39.6	38.3
	E2	37.8	36.4
	E3	35.0	33.6

TRIME measurements

SW of soil sensors.

	1 = 2.3
	2 = 2.3
	3 = 2.3

N of soil sensors.

	4 = 3.7
	5 = 3.8
	6 = 3.8

SE of soil sensors.

	7 = 2.9
	8 = 2.9
	9 = 3.0

227: LOG, Site south, Sat 22-Apr-2006 16:00:35 PDT, Guy wire tensions
With a light breeze from the south, tensions at the south tower were: 

Guy     Orientation
 #     N     SE    SW
 1    330   375   400
 2    275   320   300
 3    290   375   340
 4    280   325   300
226: krypton, Site south, Sat 22-Apr-2006 15:03:35 PDT, 5m, 30m Kryptons cleaned
22Apr06, ~12:00PDT

Cleaned with methyl alcohol

217: soil, Site south, Thu 20-Apr-2006 19:16:19 PDT, Trime Measurement, Gravimetric
20Apr06, 18:00-18:20, South

18:0PDT soil sample taken
Some roots in the 0-3cm sample, and pebbles in both.

0-3cm	164.0-79.5=85.5gm
	c1	26.3	26.1 = 0.2gm
	c2	26.5	26.3 = 0.2
	c3	23.6	23.4 = 0.2

3-6cm	160.9-44.9=116gm
	b1	28.6	28.0 = 0.6
	b2	28.6	28.0 = 0.6
	b3	35.2	34.5 = 0.7

11:40PDT Trime:
Sample	%Vol	TDR
1	2.9	64
2	2.9	64
3	6.4	62
4	6.5	62
5	6.9	46
6	7.0	46

All 3 sets of 2 readings had good penetration at 45-deg angle in the vicinity
of the station soil sensors.
Measurements 1,2 were in a very dry, loose open area.
#2,3 were in an open area under an old dead grass (clipped off to ground),
or and almost 'licen' patch.
#5,6 were taken near and towards a clump of grass under a sage bush.

216: radiation, Site south, Thu 20-Apr-2006 19:13:27 PDT, Rad.cleaned
20Apr06, 17:55

Cleaned all domes.  No real evidence of dirt on tissues or visually on domes.

198: sonic, Site south, Sat 15-Apr-2006 19:09:54 PDT, Boom angles
Poulos and Militzer shot boom angles with the new Theodolite. The leveling bubble was noted as slightly misaligned but tolerably so. John will pursue adjustment tomorrow. Shots were taken from the road west of the tower at ~ 60 m distance. Leveling was completed using a method of dual confirmation of boom parallel siting, theodolite leveling and predicted and actual sun targeting reading. The original records for the table below are noted in Poulos' green record notebook.

Sonic   Solar    Solar  Solar    Solar      Actual    Resulting   Corrected
Height  time of  Elev   Azim      Azim      Azim      Boom angle  Boom angle
        Record   stor   stor    stor d m    Meas'd    Meas        (cardinal)
------  ------   -----  ------  ----------  --------  ----------  ----------
05 m     1625    54.75  254.78  254d 46.8m  255d 05m   91d 11m    90d 52.8m
10 m     1810    75.57  271.83  271d 49.8m  272d 00m   90d 02m    89d 49.8m
15 m     1625    54.75  254.78  254d 46.8m  255d 05m   91d 11m    90d 52.8m
20 m     1740    69.56  267.32  267d 19.2m  267m 37m   91d 37m    91d 19.2m
25 m     1625    54.75  254.78  254d 46.8m  255d 05m   91d 11m    90d 52.8m
30 m     1740    69.56  267.32  267d 19.2m  267m 37m   91d 37m    91d 19.2m
184: soil, Site south, Thu 13-Apr-2006 13:16:42 PDT, Updates on Echo/Qsoil negative at south
These notes will cover a few days...

Now that we can plot the data more clearly, the soil moisture at south has been reporting negative values the entire project.
Setup spare logger and 'south' soil with spare Echo in base.  This showed roughly 350+ mV in extremely dry soil.   When I tried the probe in soil sample from central it rose to ~1V and consistent with what that station shows.
At south I tried the spare probe with the station logger with the data scope again.  This time I was more prepared to record the signal properly which was not the case yesterday due to wind/connection issues.  Today both probes showed about 380mV or roughly 2.5% per the 'standard' calibration.   I also reproduced this with the spare logger.   I then dumped some water in the 'sample soil' and watched the spare probe go up to ~700mV roughly = 16%, but this seems very low for what was actually in the box.   I also reproduced this off-line with the station logger.

CONCLUSION: my guess is the station Echo probe / logger are working fine.  The soil there appears to be different perhaps because of excess sand or salts (or less than at central).  As a result the standard calibration is probably not correct (refer to Decagon's literature which mentions this issue) don't believe the reported values!
We will bring sample containers of representative dirt from all 3 sites back to Boulder for later evaluation.

Checked Echo/logger at south
E2 excitation - per data scope is getting up to about 2500mV, and reading is at ~50mV.  Appears to confirm the logger excitation is going high enough.
Soil Sample - taken at both south and central.  Central was definitely much more moist, but to the untrained eye the soil itself appeared similar.
Spare Echo probe comparison - The signal from both the installed sensor and the spare inserted in the 'sample soil bucket' were the same: roughly 25mV raw and -40% per the logger message.  This reported value seemed a bit low per the above, so will try testing in the base tomorrow.

Rough comparison of Echo Probes between south and central per meter.
Wiring at south was OK, although the input wire on logger 11H channel was a bit loose and needed a slight tightening.
South:		139mV at highest
Central:	300-350mV at highest
Excitation is supposed to be 2500mV and the readings roughly 10-40% of that per the manual.  The slow meter was unable to grab a proper 'max' of the excitation voltage on E2 of the logger.

164: radiation, Site south, Sun 09-Apr-2006 12:42:28 PDT, Radiometers 'Cleaned'
10:00PDT 4/9/06 Cleaned radiometer domes.
These were all in good shape and did not appear to have any 'dirt' on them.
Note: slightly adjusted Rnet level
163: soil, Site south, Sun 09-Apr-2006 12:35:29 PDT, gravimetric soil moisture
Took soil sample at south at ~10:20 PDT, April 9
Taken in 'bare soil' similar to where probes appear to be.  There were some small pebbles in the batch; removed before weigh-n-bake

0-3 cm
Total weight = 157.1 - 45.0 gm = 112.1 gm
	wet	dry
d1	26.5	26.0
d2	26.0	25.5
d3	31.1	30.5

3-6 cm
Total weight = 182.2 - 79.5 = 102.7 gm
	wet	dry
c1	25.7	24.8
c2	31.0	29.9
c3	26.7	25.9

146: radiation, Site south, Tue 04-Apr-2006 08:51:55 PDT, loggers at West and South
I changed the transmit power of the ZigBee units at west and south. They are now down to 10 dBm verses 18dBm. 
144: radiation, Site south, Mon 03-Apr-2006 15:50:26 PDT, Intermittent logger data
There is once again intermittent logger data showing up in the netcdf files and
Splus plots.  However, I have scanned an entire raw data file for a time when 
this occurs and see not missing data.
143: wireless-network, Site south, Mon 03-Apr-2006 15:42:17 PDT, replaced MaxStream radio in logger at south
REplaced 316 MHz MaxStream radio in logger with 2.4 GHz Zigbee radio.

Note that the intermittent logger data at south and west have returned.  We 
suspect a software problem at the base.
126: soil, Site south, Thu 30-Mar-2006 11:04:49 PST, Qsoil negative at south
Qsoil reads about -2.5% at south.
96: radiation, Site south, Thu 23-Mar-2006 10:38:56 PST, Maintenance at south yesterday
Steve and Tom visited south tower Mar 22 before noon for routine maintenance.
Measured guywire tensions, all above 300 lbs:

1	400	370	410
2	360	325	320
3	400	320	340
4	340	330	315

Steve cleaned radiometer domes and KH2O windows with methanol.
He wiped off the net radiometer domes.
All radiometers appeared fairly clean prior to cleaning.

We crashed NIDS by running rserial locally.  This is a known
bug.  Should have run through ASTER (ssh aster@server)

63: LOG, Site south, Thu 16-Mar-2006 14:58:19 PST, South visit
- Replaced faulty 15m TRH (704) with (tested and working) spare 008.  This one
is working.
- Recleaned krypton which again lowered voltage from 0x5a00 to 0x4f00!  We'll
need to get new, fresh water and swabs and try again.
- Took soil sample (see comment #60)

55: TRH, Site south, Wed 15-Mar-2006 14:33:46 PST, Service visit to South tower
March 15, ~9.30 am - ~10:30 am

Steve and Tom visited South tower this morning.
Replaced blown fuse on NIDS serial channel 8 - 15m TRH.  Replaced fuse but Stevefound the fan was not working - probably blew the fuse again.  (Turned off 
station to replace fuse).  Replaced TRH with s/n 702(?) and replaced fuse
again - fan is working.  However, after return to trailer, getting bad data
from TRH:  both T and RH reading > 200 and s/n reporting as 704.

Steve cleaned KH2O at 30 m (~9:55 am) and 5 m (~10:30 am).
Thus on tower from ~9:45 am until 10:40 am PST.

Apparently, cleaning the KH2O's decreased the voltage at 30 m from 1.5 V to 
1.3 V!

52: TRH, Site south, Tue 14-Mar-2006 15:54:05 PST, 15 m TRH failed at south site
15m TRH failed at south site this morning at 8:20 am local time, March 14.
It had also been bad (210 deg = overrange?) for some previous period, at least
since 00Z 3/12.

Brad and Tom visited site ~3 pm and recycled power to no avail.

49: , Site south, Tue 14-Mar-2006 09:25:28 PST, South service visit
Brad and Tom visited south site March 14, after lunch.
Replaced two of three batteries;  John and Brad returned in the evening to
replace the third (waited until had field console to shut down NIDS).
44: LOG, Site south, Sun 12-Mar-2006 16:44:19 PST, south visit
visited south ~3:30-4:15:

guy wire tensions measured.

TRIME: 4.0/2.0/2.5%

Radiometers cleaned first thing.

Installed beacon switch onto aux1.

27: sonic, Site south, Sun 05-Mar-2006 16:57:44 PST, south 30m now serialized
With the serializer swap to CSAT 1117 mentioned earlier, we had a sonic that
could go to 30m.s.  We have just installed it.  That means, for the first time
for TREX, all standard sensors finally are working!  (Only 5 days and 1 1/4
IOPs late.)

Naturally, the network has started being flakey just to spite us.  My guess
is that the wind is moving the antennae around (though the TPOP looks quite
stable at the base).  The usb drives seem to be working through this.  I've
just played around with TPOP channels.  Gordon had left it as channel 11, but
I'm now running on 8 (after trying 2,3, and 5).

25: sonic, Site south, Sun 05-Mar-2006 13:28:56 PST, Serializer WARNING!
Since our now 2 spare sonics consisted of a dead sonic with a serializer and
a good sonic without a serializer (and we need a serializer still at 30m.s),
We just pulled the serializer out of 672 and installed it in 1117.  Note that
this breaks the sequence that sonics with S/N <700 had serializers and newer
ones didn't.  Someone will have to make a decision later as to whether to 
swap this serializer back.

15: data-system, Site south, Tue 28-Feb-2006 19:31:30 PST, usb drive install
The usb drive had never been installed at South.  Started it about 19:00 -- our
first data from South :(.

7: wireless-network, Site south, Mon 27-Feb-2006 20:03:42 PST, South network task
The Bulgin connector on South's etherant was damaged during installation, but
we don't want to have to strip it from all the way up the tower.
At the moment, our workaround was to install another etherant at the bottom
of the tower (using rope!), since it still has good line-of-sight to West's
AP24.  This was using another, short, cable.

Brad has rebuilt the pins on the cable to the original etherant, but was 
unable to finish installing the housing and backshell due to intense blowing
dust yesterday afternoon.  On our next visit, he'll finish this task and we
can get back to a normal configuration.  The top etherant will have to be
reconfigured since it is set now to talk to the TPOP (.12 subnet), rather than
the AP24 external antenna (.14 subnet).